OpenStudy (raphaelll):

The number 32 to the power of 1 over 5 is equal to 2. Explain why this is so, what does it mean when a number has a fractional index? Write this number without using an index

OpenStudy (danjs):

When you have a power of a fraction, say 1/n, it is the same as this: \[a ^{1/n} = \sqrt[n]{a}\]

OpenStudy (danjs):

" a to the power of 1/n, is the same as the nth root of a"

OpenStudy (danjs):

what does that tell you about 32 to the 1/5th power?

OpenStudy (raphaelll):

\[\sqrt[5]{32}\] ?

OpenStudy (danjs):

correct, that is the same as, "what number multiplied to itself 5 times is 32" n x n x n x n x n = 32?

OpenStudy (raphaelll):

which is 6.4 right?

OpenStudy (asnaseer):

@raphaelll - you have worked out a number that, when multiplied by 5, will give 32. i.e. \(6.4\times5=32\) what @DanJS was asking you is what number, when multiplied by ITSELF 5 times, will give you 32. e.g. 3 multiplied by itself 5 times will give you: \(3\times3\times3\times3\times3=243\)

OpenStudy (danjs):

Right, so this discussion was an explanation to answer the question.

OpenStudy (danjs):

2 to the 5th power = 5 *5*5*5*5 = 32 and \[32 ^{1/5} = \sqrt[5]{32} = 2\]

OpenStudy (raphaelll):

Oh I see! That makes so much sense, thanks guys